The perimeter of a rectangle is 238 cm. The length is 47 cm greater than the width. Find the length and width.
Answer by TimothyLamb(4379)

x = length in cm
y = width in cm
---
2x + 2y = 238
x = y + 47
---
put the system of linear equations into standard form
---
2x + 2y = 238
x - y = 47
---
copy and paste the above standard form linear equations in to this solver:
http://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 83
y = 36
---
Free algebra tutoring live chat:
http://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
http://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
http://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
http://sooeet.com/math/system-of-linear-equations-solver.php
---